OpenStudy (anonymous):

Factor completely: 2z(4x-y)-z(4x-y)

OpenStudy (tkhunny):

Can you do this one: 2wr - wr

OpenStudy (anonymous):

2w^2r^2

OpenStudy (anonymous):

??

OpenStudy (anonymous):

I'm having troubles

OpenStudy (tkhunny):

No good. 2wr - wr Common w w(2r - r) Common r wr(2 - 1) = wr(1) = wr How did you get squares?

OpenStudy (anonymous):

but I thought it would be (2z-z)(4x-y)

OpenStudy (anonymous):

a bit confusing.

OpenStudy (tkhunny):

Look for common factors 2z(4x-y)-z(4x-y) Common z z[2(4x-y)-(4x-y)] Common (4x-y) z(4x-y)[2 - 1] = z(4x-y)[1] = z(4x-y) Yours is fine. (2z-z)(4x-y) You just need to find the z and you'll nearly be done.

OpenStudy (anonymous):

thank you for explaining :)